OpenStudy (anonymous):

what is 3x-9y=0 -x+3y=-3

OpenStudy (anonymous):

fast

OpenStudy (timmyk14):

Post the full question please

OpenStudy (anonymous):

it is the full question

OpenStudy (anonymous):

I assume you need us to solve the system? By what? Elimination? Substitution?

OpenStudy (anonymous):

the rest is how many soloutions are there to the following system of equations 3x-9y=0 -x+3y=-3

OpenStudy (anonymous):

the choices are 2 0 infinitley many 1

OpenStudy (timmyk14):

@BunnyBree just explained what is you have to do to solve this problem. Do you understand how to solve these equation? If not we can help you, and guide you to the correct answer.

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

i need the way to solve it with the answer

OpenStudy (anonymous):

fast plz i still got 10 more questions to do